OpenStudy (anonymous):

If f(x)=x + 2 and g(x)=x^2 - 4x, find: f(g(-3))

razor99 (razor99):

Thats functions right

OpenStudy (anonymous):

Yes

OpenStudy (ranga):

evaluate g(-3) first amd then put that in the place of x in f(x) and evaluate the function.

OpenStudy (anonymous):

So this would be correct?\[-3^2-4(-3)\]

OpenStudy (ranga):

parenthesis around (-3)

OpenStudy (whpalmer4):

Not quite. \[g(x) = x^2-4x\]\[x=-3\]\[g(-3) = (-3)^2-4(-3) =\] then whatever that works out to be, plug it in as \(x\) in \(f(x) = x+2\)

OpenStudy (anonymous):

Then the final answer to that would be 5 after I take the 3 and plug it in for f(x)?

OpenStudy (whpalmer4):

need to check your arithmetic.

OpenStudy (whpalmer4):

what is -3 * -3?

OpenStudy (anonymous):

-9?

OpenStudy (whpalmer4):

then what is 3 * -3?

OpenStudy (anonymous):

No, 9.

OpenStudy (whpalmer4):

yeah, so that's \[(-3)^2-4(-3) = 9-(-12) = \]

OpenStudy (anonymous):

21

OpenStudy (whpalmer4):

right! so \(g(-3) = 21\) now what does \(f(21)=\)

OpenStudy (anonymous):

23 because the f(21) would be 21 + 2

OpenStudy (whpalmer4):

excellent! \[f(g(-3)) = f(21) = 23\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (whpalmer4):

it can be a bit confusing at first, because the \(x\) gets used over and over again and isn't necessarily the same value. think of the f(x) = <something> as a recipe...